The dielectric strength of air is 3.0 × 10^6 V/m. A parallel-plate air-capacitor has area 20 cm^2 and plate separation 0.10 mm. Find the maximum rms voltage of an AC source which can be safely connected to this capacitor.

Grade:upto college level

1 Answers

Aditi Chauhan
9 years ago
Sol. E = 3 * 10^6 V/m, A = 20 cm^2, d = 0.1 mm Potential diff. across the capacitor = Ed = 3 * 10^6 * 0.1 * 10^-3 = 300 V Max. rms Voltage = V/√2 = 300/√2 = 212 V